I need help with this problem please..
Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
-3x + 4y = 1, -6x + 6y = -3

Answer by mananth(16949)

-3 x + 4 y = 1 .............1
-6 x + 6 y = -3 .............2
Eliminate y
multiply (1)by -3
Multiply (2) by 2
9 x -12 y = -3
-12 x 12 y = -6
Add the two equations
-3 x = -9
/ -3
x = 3
plug value of x in (1)
-3 x + 4 y = 1
-9 + 4 y = 1
4 y = 1 + 9
4 y = 10
y = 2.5